how do you arrange these physics equations
t=2pi sqrt((r^3)/(gm)) solve for m
a/b + c= a/d solve for b
d = vt + (1/2)at2 solve for v

Found 2 solutions by texttutoring, solver91311:
Answer by texttutoring(324)

t=2pi sqrt((r^3)/(gm))
Divide both sides by 2Pi

T/2Pi = sqrt(r^3/gm)

Square both sides:

(T/2Pi)^2 = r^3/gm

Multiply both sides by m:

m*(T/2Pi)^2 = r^3/g

Divide both sides by (T/2Pi)^2:
m = (r^3/g)(2Pi/T)^2


Next question: a/b + c= a/d
a/b=a/d -c
a = (a/d-c)b
a/(a/d - c) = b

Last question: d = vt + (1/2)at^2
d = vt + (1/2)at^2
d - (1/2)at^2 = vt

Divide both sides by t:

d/t - (1/2)at = v

Answer by solver91311(24713)





Swap sides:



Multiply both sides by :



Square both sides:



Multiply both sides by :



Take the reciprocal of both sides:



Use similar techniques to solve the other two problems.


John

My calculator said it, I believe it, that settles it